OpenStudy (anonymous):

there's 10 white golf balls 6 stripped golf balls the golfer wants to add 112 golf balls to the bag but wants the ratio of white to stripped golf balls to stay the same. how many of each should he add?

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@AriPotta @Ammarah @Awstinf

OpenStudy (anonymous):

@bloopman @Blu-Girl @ButterflyHope @biancanatel @bkonch @bugzkilla @Clalgee @caozeyuan @coolsday @ComeAlongP0nd @Candycane @confusegirllll @dan815 @DullJackel09 @DrPepperx3 @danish071996 @dabeezy @dahlinal @ElisaNeedsHelp @ememlove @elleblythe

OpenStudy (anonymous):

@emo__gamer101 @Erykahbabe @fortheloveofwolves @fortheloveofwolves @fustration @FibonacciChick666 @ganeshie8 @Gixxer_J @gdon384 @gibblegriggs @geerky42 @Haley1654 @HugoHuynh @iambatman @jim_thompson5910 @Jhannybean @Johnbc @JackleBee @juliamae10 @jschwa01 @kropot72 @KittyT @kimberlyciancaglione @KonaFirestar @khusker @Korey_Minter @liz_mustang @lily1021 @lovetime121 @lift.mosh.repeat_ @lift.mosh.repeat_ @lindsay22813 @love.dove

OpenStudy (anonymous):

I am of no help whatsoever

OpenStudy (anonymous):

Sorry, I totally hate math with a burning passion

OpenStudy (anonymous):

@magepker728 @Mandre @MilkNCookies @missmysterious @magy33 @Mylove! @nikato @NoelGreco @naruto21 @OrangeBlood @One_3234 @perl @Polkapen @prettyboy937 @prettyboy937 @paghesp @quickstudent @Qawi @razor99 @RyGuy @roy.g.biv @RJaza @skullpatrol @sammixboo @ShadowLegendX @swissgirl @shadow22 @SydtheKid913 @timo86m

OpenStudy (anonymous):

@TheSmartOne @TheSecretsToday @Turner @tayromo @Tedimariah @vexus @VJT @Vickiesweet @vanesilva496 @Werewolfprincess @willet @whead1 @Yuba @yoitsren @Yocy @z053sme @ziggypiggy22

TheSmartOne (thesmartone):

Please do not mass tag.

TheSmartOne (thesmartone):

That is against the CoC.

OpenStudy (anonymous):

\[\frac{ White..golf }{ total..golf } = \frac{ white...golf }{ total...golf}\]

OpenStudy (anonymous):

there's 10 white golf balls 6 stripped golf balls the golfer wants to add 112 golf balls to the bag but wants the ratio of white to stripped golf balls to stay the same. how many of each should he add? 10 wb ------ = 16 6 sb 112-16= 96 add hold on

OpenStudy (anonymous):

per every 5 wb you add you must add 3 sb right? according to the ratio

OpenStudy (anonymous):

The ratio of white golf balls before over total and white golf balls after over total after\[\frac{ 10 }{ 16 } = \frac{ X }{ 112}\]

OpenStudy (anonymous):

for striped set up a proportion of striped balls to striped/white balls (6/16) and set it equal to x/total balls (x/112)

OpenStudy (anonymous):

john bc is the proportion to solve for white balls and the one I wrote will work for the striped balls. Those are the only two you need

OpenStudy (anonymous):

same as saying per every 1 wb add .625 sb. ANd so Totalballs= 16 +(wb*x)+(sb*x) 1 wb is just 1 and one sb is .625. Total needed is 112. so plug in 112=16+(x)+(.635*x) solve for x

OpenStudy (anonymous):

58.7155963 wb and 58.7155963*.625 = 37 sb you get 96 or 97 depending on how you round + the 16 already in the bag is 112 RIGHT?

OpenStudy (anonymous):

in other words add 58 or 59 white balls. THEN SB equal number_you_chose * .625 .625 is the same ratio

OpenStudy (anonymous):

@AccessDenied @AriPotta @Ammarah @Awstinf @bloopman @Blu-Girl @ButterflyHope @biancanatel @bkonch @bugzkilla @Clalgee @caozeyuan @coolsday @ComeAlongP0nd @Candycane @confusegirllll @dan815 @DullJackel09 @DrPepperx3 @danish071996 @dabeezy @dahlinal @ElisaNeedsHelp @ememlove @elleblythe @emo__gamer101 @Erykahbabe @fortheloveofwolves @fortheloveofwolves @fustration @TheSmartOne @TheSecretsToday @Turner @tayromo @Tedimariah @vexus @VJT @Vickiesweet @vanesilva496 @Werewolfprincess @willet @whead1 @Yuba @yoitsren @Yocy @z053sme @ziggypiggy22 @FibonacciChick666 @ganeshie8 @Gixxer_J @gdon384 @gibblegriggs @geerky42 @Haley1654 @HugoHuynh @iambatman @jim_thompson5910 @Jhannybean @Johnbc @JackleBee @juliamae10 @jschwa01 @kropot72 @KittyT @kimberlyciancaglione @KonaFirestar @khusker @Korey_Minter @liz_mustang @lily1021 @lovetime121 @lift.mosh.repeat_ @lift.mosh.repeat_ @lindsay22813 @love.dove @magepker728 @Mandre @MilkNCookies @missmysterious @magy33 @Mylove! @nikato @NoelGreco @naruto21 @OrangeBlood @One_3234 @perl @Polkapen @prettyboy937 @prettyboy937 @paghesp @quickstudent @Qawi @razor99 @RyGuy @roy.g.biv @RJaza @skullpatrol @sammixboo @ShadowLegendX @swissgirl @shadow22 @SydtheKid913 @timo86m

OpenStudy (anonymous):

@sassydiva321 mines is by far best explanation .

TheSmartOne (thesmartone):

Stop mass tagging that is against the \(\Huge\color{red}{Code~of~Conduct}\)

OpenStudy (anonymous):

you are no fun "Smartone"

sammixboo (sammixboo):

Yes, I believe @timo86m gave a pretty good explination :) Spammy = No no Js :)

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

There are worst things going on than asking for math help. If you cant help or dont know then just ignore and move on to help someone you can.

OpenStudy (anonymous):

\[LOL\]

OpenStudy (anonymous):

add the stripped golf balls to the white golf balls which would be ____ then divide that number to 112

OpenStudy (anonymous):

@timo86m do you just monitor people on here???

OpenStudy (anonymous):

and thank you @vanesilva496 that helped a lot

OpenStudy (anonymous):

your welcome :D i just give out simple equations while everyone is doing very advanced equations. I really do hope my response helped you out with your question

OpenStudy (anonymous):

\[Thank!you!so!much!VANSILVA\]

OpenStudy (anonymous):

\[Thank!you!so!much!TIMO\]

OpenStudy (anonymous):

@timo86m @vanesilva496